# Online Difficulty Filtering for Reasoning Oriented Reinforcement Learning

### v1

@@ Abstract
Reasoning-Oriented Reinforcement Learning (RORL) enhances the reasoning ability of Large Language Models (LLMs). However, due to the sparsity of rewards in RORL, effective training is highly dependent on the selection of problems of appropriate difficulty. Although curriculum learning attempts to address this by adjusting difficulty, it often relies on static schedules, and even recent online filtering methods lack theoretical grounding and a systematic understanding of their effectiveness. In this work, we theoretically and empirically show that curating the batch with the problems that the training model achieves intermediate accuracy on the fly can maximize the effectiveness of RORL training, namely balanced online difficulty filtering. We first derive that the lower bound of the KL divergence between the initial and the optimal policy can be expressed with the variance of the sampled accuracy. Building on those insights, we show that balanced filtering can maximize the lower bound, leading to better performance. Experimental results across five challenging math reasoning benchmarks show that balanced online filtering yields an additional 10% in AIME and 4% improvements in average over plain GRPO. Moreover, further analysis shows the gains in sample efficiency and training time efficiency, exceeding the maximum reward of plain GRPO within 60% training time and the volume of the training set.
@@ 1 Introduction
Reinforcement Learning (RL) has become a key training paradigm for training large language models (LLMs) specialized in reasoning tasks, exemplified by OpenAI o1 (OpenAI et al., 2024) and DeepSeek-R1 (Guo et al., 2025). These models utilize Reasoning-Oriented Reinforcement Learning (RORL), where verifiable rewards like correctness in mathematical or logical problems serve as the primary supervision signal (Lambert et al., 2024).
@@
As RORL increasingly targets high-complexity reasoning tasks, designing effective learning dynamics becomes crucial to help models progressively acquire the necessary capabilities. Effective learning has long been studied in the education domain, where theories such as the Zone of Proximal Development (ZPD) (Cole, 1978; Tzannetos et al., 2023) emphasize that learning is most efficient when tasks are neither too easy nor too hard, but instead fall within a learner’s optimal challenge zone. This has motivated a variety of strategies in language modeling, from curriculum learning that introduces harder problems progressively (Team et al., 2025), to difficulty-aware data curation that selects or filters examples based on estimated pass rates or diversity (Muennighoff et al., 2025; Ye et al., 2025). Online filtering methods further explore this idea by dynamically adjusting the training data to match the current ability of the model (Cui et al., 2025). However, while previous work demonstrate the empirical effectiveness of such techniques, they often lack a detailed analysis of why or when certain difficulty distributions yield better learning outcomes.
In this work, we conduct extensive experiments and provide theoretical analysis to understand how and why difficulty filtering improves learning in RORL. We start by deriving that the lower bound of the KL divergence between the learned policy and the optimal policy is proportional to the sample accuracy, and this divergence is theoretically maximized when the pass rate is around 0.5. Based on this insight, we focus on balanced online difficulty filtering (Figure 1), which maintains a range of problem difficulties centered around the current ability of the model. This approach improves learning efficiency by keeping training examples within the predefined difficulty range, where each batch maximizes its expected learning signal. In practical implementation, we avoid the instability caused by prior methods that naively discard overly easy or hard examples (Cui et al., 2025; Meng et al., 2025). Instead, we replace filtered-out samples with others using parallel sampling, ensuring consistent batch sizes and time-efficient training.
Experiments on five challenging mathematical reasoning benchmarks (Hendrycks et al., 2021; Li et al., 2024; Lewkowycz et al., 2022; He et al., 2024) show that this online difficulty filtering significantly outperforms both non-curriculum and offline curriculum baselines, highlighted by exceeding plain GRPO by 10% points in AIME and offline filtering by 4.2% points in average. We find that balanced filtering—removing both easy and hard problems—improves sample efficiency and final performance, while skewed filtering leads to suboptimal learning. Moreover, the method adapts dynamically as the model improves, providing similar benefits to curriculum learning while avoiding the limitations of static schedules. Our findings highlight the importance of dynamic, balanced difficulty control in reinforcement learning, demonstrating a principled and efficient method for RORL.
@@
Figure 1: Balanced online difficulty filtering for maximizing the effectiveness of GRPO. With $G$ rollouts for each prompt $\mathbf{x}$, we measure the pass rate $p(\mathbf{x})$ as the average accuracy and filter them by predefined thresholds: e.g., $0.25\leq p(\mathbf{x})\leq 0.75$ in this case. We recursively stack filtered prompts until the train batch size meets the fixed size $N$. We elaborate on the asynchronous implementation in Appendix A.
2 Related Works
Reasoning-oriented reinforcement learning.
Recent advancements demonstrate significant reasoning improvements in LLMs through RL (Havrilla et al., 2024; OpenAI et al., 2024; Lambert et al., 2024; Guo et al., 2025; OLMo et al., 2025; Kumar et al., 2025). OpenAI o1 (OpenAI et al., 2024) initially reported that increasing the compute during RL training and inference improves reasoning performance. DeepSeek R1 (Guo et al., 2025) further found that, in RORL with verifiable rewards, longer responses correlate with better reasoning. Concurrent studies (Team et al., 2025; Hou et al., 2025; Luo et al., 2025) employed algorithms, such as GRPO (Shao et al., 2024) or RLOO (Ahmadian et al., 2024), relying on advantage estimation via multiple response samples rather than PPO-like value networks. Hou et al. (2025) further found that training efficiency improved with increased sampling in RLOO, invoking the need for more sample-efficient training strategies in reasoning-oriented RL.
Difficulty-based curriculum learning.
Curriculum learning has been widely adopted in fine-tuning LLMs to improve training efficiency (Lee et al., 2024; Naïr et al., 2024; Team et al., 2025; Cui et al., 2025). Static curricula, i.e., offline data curation with a predetermined task difficulty, have been effective in multiple domains: instruction-tuning (Lee et al., 2024) and coding (Naïr et al., 2024; Team et al., 2025; Li et al., 2025) to name a few. In RORL, Team et al. (2025) employs a static difficulty-based curriculum, assigning tasks at fixed difficulty levels to ensure efficient progression. Similarly, Li et al. (2025) selects a high-impact subset of training data based on a “learning impact measure”. Meantime, adaptive curricula dynamically adjust task difficulty based on the learners’ progress, addressing the limitations of static curricula (Florensa et al., 2018; Cui et al., 2025). Specifically, Cui et al. (2025) applied adaptive filtering in reasoning and reported an empirical advantage in reducing reward variance. However, Meng et al. (2025) observed that such dynamic exclusion of examples may destabilize training, as it causes fluctuations in the effective batch size.
@@ 3 Preliminaries
Given the training policy $\pi_{\theta}$ initialized from the reference policy $\pi_{\mathrm{init}}$, reinforcement learning (RL) in language model environment optimizes $\pi_{\theta}$ to maximize the reward assessed by the reward function $r$ (Christiano et al., 2017; Ziegler et al., 2020) through:
$\max_{\theta}\mathbb{E}_{\mathbf{y}\sim\pi_{\theta}(\cdot|\mathbf{x})}[r(\mathbf{x},\mathbf{y})-\beta\mathbb{D}_{\mathrm{KL}}\left(\pi_{\theta}(\mathbf{y}|\mathbf{x})|\pi_{\mathrm{init}}(\mathbf{y}|\mathbf{x})\right)],$ (1)
@@
$\pi^{*}(\mathbf{y}|\mathbf{x})=Z(\mathbf{x})\pi_{\mathrm{init}}(\mathbf{y}|\mathbf{x})\exp\left(\frac{1}{\beta}r(\mathbf{x},\mathbf{y})\right),$ (3)
@@
Unlike PPO, recent works propose excluding parameterized value models (Ahmadian et al., 2024; Kazemnejad et al., 2024; Wu et al., 2024), including group relative policy optimization (Shao et al., 2024, GRPO).
GRPO leverages the PPO-style clipped surrogate objective but calculates the policy gradient by weighting the log-likelihood of each trajectory with its group-based advantage, thus removing the need for a critic (Vojnovic and Yun, 2025; Mroueh, 2025). For each prompt, $G$ parallel responses are sampled and their reward $r_{i}$ is used to calculate the advantage $\hat{A}_{i}$:
@@
where $\mathrm{mean}(\cdot)$ and $\mathrm{std}(\cdot)$ are the average and standard deviation of the input values. The effectiveness of GRPO is especially highlighted in the tasks with verifiable reward stipulated through the binary reward functions (Lambert et al., 2024; Guo et al., 2025; Wei et al., 2025):
@@ 4 Learnability in GRPO and Online Difficulty Filtering
In this section, we analyze the learnability of the prompt in RL with language model environments under binary rewards. We show that prompts that are either too easy or too hard yield no learning signal (§4.2), while intermediate ones—characterized by high reward variance—maximize the gradient information (§4.3). Building on these insights, we propose a balanced online difficulty filtering4.4 and §4.5) to optimize GRPO training.
4.1 Background: Prompt-level learnability
We begin by recalling the definition of the optimal value function and the partition function in the soft RL setting (Schulman et al., 2018; Richemond et al., 2024):
$V^{*}(\mathbf{x}):=\beta\log\mathbb{E}_{\mathbf{y}\sim\pi_{\mathrm{init}}(\cdot|\mathbf{x})}\left[\exp\left(\frac{1}{\beta}r\left(\mathbf{x},\mathbf{y}\right)\right)\right]~{}\mathrm{and}~{}Z(\mathbf{x})=\exp\left(\frac{1}{\beta}V^{*}(\mathbf{x})\right).$ (6)
Using $V^{*}(\mathbf{x})$ in Equation (3), the log ratio between the initial policy $\pi_{\mathrm{init}}$ and the optimal policy $\pi^{*}$ can be expressed as:
@@ 4.1 Background: Prompt-level learnability
$\mathbb{E}_{\mathbf{y}\sim\pi_{\mathrm{init}}(\cdot|\mathbf{x})}\left[\log\frac{\pi^{*}(\mathbf{y}|\mathbf{x})}{\pi_{\mathrm{init}}(\mathbf{y}|\mathbf{x})}\right]=\frac{1}{\beta}\mathbb{E}_{\mathbf{y}\sim\pi_{\mathrm{init}}(\cdot|\mathbf{x})}\left[r(\mathbf{x},\mathbf{y})\right]-\frac{1}{\beta}V^{*}(\mathbf{x}),$ (8)
where the right-hand side (RHS) represents a soft-RL variant of the advantage function scaled by $\beta^{-1}$ (Haarnoja et al., 2017; Schulman et al., 2018), as $\mathbb{E}_{\pi_{\mathrm{init}}}\left[r(\mathbf{x},\mathbf{y})\right]$ can be interpreted as Q-function. And the left-hand side (LHS) corresponds to the negative reverse KL divergence between $\pi_{\mathrm{init}}$ and $\pi^{*}$ (Rafailov et al., 2024):
$\mathbb{D}_{\mathrm{KL}}\left(\pi_{\mathrm{init}}(\mathbf{y}|\mathbf{x})|\pi^{*}(\mathbf{y}\,\|\,\mathbf{x})\right)=-\mathbb{E}_{\mathbf{y}\sim\pi_{\mathrm{init}}(\cdot|\mathbf{x})}\left[\log\frac{\pi^{*}(\mathbf{y}|\mathbf{x})}{\pi_{\mathrm{init}}(\mathbf{y}|\mathbf{x})}\right].$ (9)
@@
For the binary reward $r_{\mathrm{acc}}$ in Equation (5), the reward distribution is Bernoulli with parameter $p(\mathbf{x})$ for prompt $\mathbf{x}$, policy $\pi$, and $\mathbf{y}\sim\pi(\cdot|\mathbf{x})$, which we refer to as “pass rate”:
@@
    Absolute-hard prompts ($\mathbf{x}_{\mathrm{Hard}},~{}p(\mathbf{x}_{\mathrm{Hard}})=0$)
@@
    Soft-hard prompts ($\mathbf{x}_{\mathrm{hard}},p(\mathbf{x}_{\mathrm{hard}})=\epsilon$)
@@
    Intermediate prompts ($\mathbf{x}_{\mathrm{inter}},~{}\epsilon\leq p(\mathbf{x}_{\mathrm{inter}})\leq 1-\epsilon$)
@@
    Soft-easy prompts ($\mathbf{x}_{\mathrm{easy}},~{}p(\mathbf{x}_{\mathrm{easy}})=1-\epsilon$)
@@
    Absolute-easy prompts ($\mathbf{x}_{\mathrm{Easy}},~{}p(\mathbf{x}_{\mathrm{Easy}})=1$)
@@
where $\epsilon$ is a small positive constant satisfying $0.50\ll\epsilon<$ 0.5. The variance is zero if and only if $p(\mathbf{x})=0$ or $p(\mathbf{x})=1$, corresponding to absolute hard and absolute easy prompts, respectively.
4.2 Case 1: Learnability in absolute prompts
For absolute prompts $\mathbf{x}_{\mathrm{Hard}}$ and $\mathbf{x}_{\mathrm{Easy}}$, both the expected reward and the state value are zero and one, respectively:
$\begin{cases}~{}~{}\mathbb{E}_{\pi_{\mathrm{init}}}\left[r_{\mathrm{acc}}(\mathbf{x},\mathbf{y})\right]=0~{}~{}\mathrm{and}~{}~{}V^{*}(\mathbf{x})=\beta\log\left(\exp\left(\frac{0}{\beta}\right)\right)=0.&\quad\text{if}~{}~{}~{}\mathbf{x}=\mathbf{x}_{\mathrm{Hard}},\\ ~{}~{}\mathbb{E}_{\pi_{\mathrm{init}}}\left[r_{\mathrm{acc}}(\mathbf{x},\mathbf{y})\right]=1~{}~{}\mathrm{and}~{}~{}V^{*}(\mathbf{x})=\beta\log\left(\exp\left(\frac{1}{\beta}\right)\right)=1.&\quad\text{if}~{}~{}~{}\mathbf{x}=\mathbf{x}_{\mathrm{Easy}}.\end{cases}$ (11)
By Equation (8), the expected log ratio between $\pi_{\mathrm{init}}$ and $\pi^{*}$ become zero, implying that $\pi_{\mathrm{init}}$ is already optimal regarding the initial model’s capability:
$\mathbb{D}_{\mathrm{KL}}\left(\pi_{\mathrm{init}}(\mathbf{y}|\mathbf{x})\,\|\,\pi^{*}(\mathbf{y}|\mathbf{x})\right)=0~{}\mathrm{when}~{}\mathbf{x}\in\left\{\mathbf{x}_{\mathrm{Hard}},\mathbf{x}_{\mathrm{Easy}}\right\}.$ (12)
Therefore, absolute-hard and absolute-easy prompts ($p(\mathbf{x})\in\left\{0,1\right\}$) do not contribute useful gradient information during RL training. For GRPO in specific, this is an intuitive result as the advantage $\hat{A}$ in GRPO naturally become zero for every rollout by Equation (4).
4.3 Case 2: Learnability in soft prompts
Next, we show that the prompts with $p(\mathbf{x})\simeq 0.5$ have the largest learnability, thereby preserving the prompts with $\epsilon\leq p(\mathbf{x})\leq 1-\epsilon$ maximize the effectiveness in the RL phase.
Reward variance as a lower bound of optimal divergence.
Regarding $r_{\mathrm{acc}}(\mathbf{x},\mathbf{y})$ with $\mathbf{y}\sim\pi_{\mathrm{init}}(\cdot|\mathrm{x})$ is Bernoulli, we can rewrite $V^{*}(\mathbf{x})$ as:
$V^{*}(\mathbf{x})=\beta\log\left((1-p(\mathbf{x}))+p(\mathbf{x})\exp\left(\frac{1}{\beta}\right)\right)$ (13)
by substituting $\mathbb{E}_{\pi_{\mathrm{init}}}\left[\exp\left(r(\mathbf{x},\mathbf{y})/\beta\right)\right]$ through a simple exponential transformation of Bernoulli distribution. With the second order Taylor expansion of $\exp\left(1/\beta\right)$ and applying it to Equation (8), we have:

### v2

@@ Abstract
Recent advances in reinforcement learning with verifiable rewards (RLVR) show that large language models enhance their reasoning abilities when trained with verifiable signals. However, due to reward sparsity, effectiveness depends heavily on selecting samples of appropriate difficulty. In this work, we present a formal analysis of online difficulty-aware filtering and establish its theoretical foundations. We show that expected policy improvement is lower-bounded by the variance of task-level success probabilities, implying that selecting tasks of intermediate difficulty maximizes learning efficiency. Building on this, we demonstrate that balanced filtering maximizes this lower bound, leading to superior performance and sample efficiency. Evaluations across multiple math reasoning benchmarks validate that balanced filtering consistently enhances convergence speed and final performance, achieving up to +12% gains in less than half the training steps of standard GRPO. By extending our analysis to various reward distributions, we provide a principled foundation for future RLVR curriculum strategies, confirmed through both theoretical analysis and extensive empirical results.
Online Difficulty Filtering for Reasoning Oriented Reinforcement Learning
Sanghwan Bae1thanks: Equal contribution. Jiwoo Hong2∗thanks: Work was done while the authors were at NAVER Cloud. Min Young Lee1 Hanbyul Kim1 JeongYeon Nam322footnotemark: 2 Donghyun Kwak1 NAVER Cloud1 KAIST AI2 TwelveLabs3 baaesh10@gmail.com, jiwoo_hong@kaist.ac.kr
@@ 1 Introduction
Reinforcement learning (RL) has become a key training paradigm for training large language models (LLMs) to further enhance their generational capabilities (Ouyang et al., 2022; Touvron et al., 2023; Yang et al., 2024a; Walsh et al., 2025), often posed as post-training. Specifically, reinforcement learning with verifiable rewards (RLVR) that have a discrete set of correct answers for domains like math reasoning is emerging as a new application of RL (OpenAI et al., 2024; Lambert et al., 2025a; Guo et al., 2025). Despite its effectiveness, the training efficiency is the main bottleneck in RL for LLMs, which recent works try to overcome either by hardware optimization (Mei et al., 2025; Noukhovitch et al., 2025) or algorithmic solutions (Lee and Lim, 2024; Ahmadian et al., 2024).
@@
Efficient learning, i.e., achieving the optimal performance with less data, has long been studied in the education domain, where theories such as the Zone of Proximal Development (Cole, 1978; Tzannetos et al., 2023, ZPD) emphasize that learning is most efficient when tasks are neither too easy nor too hard, but instead fall within a learner’s optimal challenge zone. This has motivated a variety of strategies in general language modeling (Platanios et al., 2019; Maharana and Bansal, 2022; Xie et al., 2023). When the idea of filtering the data with intermediate difficulty is applied to RLVR, it can be used for progressively introducing harder problems (Team et al., 2025) or filtering examples based on the pre-defined proxies (Muennighoff et al., 2025; Ye et al., 2025; Yang et al., 2025). Especially, by setting the proxy as the training policy’s capability, we can harness the online learning in RLVR for difficulty-aware efficient learning (Cui et al., 2025). Despite applying difficulty-aware data curation can be found in recent works for RLVR, they often lack detailed analysis on how the theory of ZPD can be linked to online reinforcement learning algorithms, i.e., theoretical foundation of online difficulty-aware filtering in RLVR.
@@
Figure 1: Balanced online difficulty filtering for maximizing the effectiveness of GRPO. With $G$ rollouts for each prompt $\mathbf{x}$, we measure the pass rate $p(\mathbf{x})$ as the average accuracy and filter them by predefined thresholds: e.g., $0.25\leq p(\mathbf{x})\leq 0.75$ in this case. We recursively stack filtered prompts until the train batch size meets the fixed size NN. We elaborate on the asynchronous implementation in Appendix B.
In this work, we establish a theoretical foundation for online difficulty-aware filtering in RLVR through identifying the sample reward variance as the lower bound of the reversed KL divergence between the initial policy and the optimal policy, thereby being an effective proxy for filtration. With a novel asynchronous sampling strategy that replaces filtered-out items with parallel rollouts and retains the batch size, we build a theoretical and empirical background a reliable filtering strategy for RLVR. Our contributions are as follows:
  1. 1.
    Theoretical assessment of prompt-level learnability3): We prove that the learning signal for the prompt, given the initial policy, can be approximated via sample reward variance (Proposition 3.1), implying why the prompts with extremely low or high pass rate should be filtered (Remarks 3.3 and 3.2).
  2. 2.
    Empirical scalability and generalizability of balanced online difficulty filtering5): With 3B and 7B models, we empirically validate the effectiveness of the balanced online difficulty filtering across five math reasoning benchmarks with varying levels, e.g., $10\%$ on AIME and $4.2\%$ in average for 3B and $12\%$ on AMC and $4.5\%$ in average for 7B.
  3. 3.
    Sample efficiency of balanced online difficulty filtering6): We show that the filtering in fact decreases the amount of training data and time in achieving the optimal performance, using less than half of gradient updates to outperform the plain GRPO with filtering.
  4. 4.
    Generalizability of sample reward variance as learnability proxy6.3): We provide a general proof that wider range of reward distributions, e.g., Gaussian (Corollary 6.2) or Multinomial (Corollary 6.3), can also enjoy the sample reward variance as the learnability proxy.
@@ 2 Preliminaries
Given the training policy $\pi_{\theta}$ initialized from the reference policy $\pi_{\mathrm{init}}$, reinforcement learning (RL) in language model environment optimizes $\pi_{\theta}$ to maximize the reward assessed by the reward function $r$ (Christiano et al., 2017; Ziegler et al., 2020):
$\max_{\theta}\mathbb{E}_{\mathbf{y}\sim\pi_{\theta}(\cdot|\mathbf{x})}\left[r(\mathbf{x},\mathbf{y})\right]-\beta\mathbb{D}_{\mathrm{KL}}\left(\pi_{\theta}\|\pi_{\mathrm{init}}\right),$ (1)
@@
$\pi^{*}(\mathbf{y}|\mathbf{x})=Z(\mathbf{x})\pi_{\mathrm{init}}(\mathbf{y}|\mathbf{x})e^{\frac{1}{\beta}r(\mathbf{x},\mathbf{y})},$ (3)
@@
Unlike PPO, recent works exclude parameterized value models (Ahmadian et al., 2024; Kazemnejad et al., 2024; Wu et al., 2024), including group relative policy optimization (Shao et al., 2024, GRPO). GRPO leverages the PPO-style clipped surrogate objective but calculates the policy gradient by weighting the log-likelihood of each trajectory with its advantage, thus removing the need for a critic (Vojnovic and Yun, 2025; Mroueh, 2025). For each prompt, $G$ sampled responses and their reward $r_{i}$ is used to calculate the advantage $\hat{A}_{i}$:
@@
where $\mathrm{mean}(\cdot)$ and $\mathrm{std}(\cdot)$ are the average and standard deviation of the input values. The effectiveness of GRPO is especially highlighted in the tasks with verifiable reward stipulated through the binary reward functions (Lambert et al., 2024; Guo et al., 2025; Wei et al., 2025b):
@@ 3 Learnability in GRPO and Online Difficulty Filtering
We study the learnability of prompts in reinforcement learning with language model environments under binary rewards. Our analysis shows that prompts that are either trivially easy or impossibly hard yield zero divergence and thus no learning signal, while intermediate prompts with higher reward variance maximize the effective gradient information. These results are formalized in Proposition 3.1, which motivates a balanced online difficulty filtering strategy (§3.33.4) for optimizing GRPO training.
3.1 Background
The optimal value function and the partition function in the soft RL setting (Schulman et al., 2018; Richemond et al., 2024) are defined as:
$V^{*}(\mathbf{x}):=\beta\log\mathbb{E}_{\mathbf{y}\sim\pi_{\mathrm{init}}(\cdot|\mathbf{x})}\left[e^{\frac{1}{\beta}r\left(\mathbf{x},\mathbf{y}\right)}\right]$ (6)
$Z(\mathbf{x})=\exp\left(\frac{1}{\beta}V^{*}(\mathbf{x})\right).$ (7)
Using $V^{*}(\mathbf{x})$ in equation 3, the log ratio between $\pi_{\mathrm{init}}$ and $\pi^{*}$ can be expressed as:
@@ 3.1 Background
$\begin{split}&\mathbb{E}_{\mathbf{y}\sim\pi_{\mathrm{init}}(\cdot|\mathbf{x})}\left[\log\frac{\pi^{*}(\mathbf{y}|\mathbf{x})}{\pi_{\mathrm{init}}(\mathbf{y}|\mathbf{x})}\right]\\ &=\frac{1}{\beta}\mathbb{E}_{\mathbf{y}\sim\pi_{\mathrm{init}}(\cdot|\mathbf{x})}\left[r(\mathbf{x},\mathbf{y})\right]-\frac{1}{\beta}V^{*}(\mathbf{x}),\end{split}$ (9)
where the right-hand side (RHS) is a soft-RL variant of the advantage function scaled by $\beta^{-1}$ (Haarnoja et al., 2017; Schulman et al., 2018), as $\mathbb{E}_{\pi_{\mathrm{init}}}\left[r(\mathbf{x},\mathbf{y})\right]$ can be interpreted as a Q-function. And the left-hand side (LHS) corresponds to the negative reverse KL divergence between $\pi_{\mathrm{init}}$ and $\pi^{*}$ (Rafailov et al., 2024):
$\begin{split}&\mathbb{D}_{\mathrm{KL}}\left(\pi_{\mathrm{init}}(\mathbf{y}|\mathbf{x})|\pi^{*}(\mathbf{y}\,\|\,\mathbf{x})\right)\\ &~~=-\mathbb{E}_{\mathbf{y}\sim\pi_{\mathrm{init}}(\cdot|\mathbf{x})}\left[\log\frac{\pi^{*}(\mathbf{y}|\mathbf{x})}{\pi_{\mathrm{init}}(\mathbf{y}|\mathbf{x})}\right].\end{split}$ (10)
@@
For the binary reward $r_{\mathrm{acc}}$, the reward distribution is Bernoulli with parameter $p(\mathbf{x})$ for prompt $\mathbf{x}$, policy $\pi$, and $\mathbf{y}\sim\pi(\cdot|\mathbf{x})$, which we refer as “pass rate”:
@@
    Absolute-hard ($\mathbf{x}_{\mathrm{Hard}},~p(\mathbf{x}_{\mathrm{Hard}})=0$)
@@
    Soft-hard ($\mathbf{x}_{\mathrm{hard}},p(\mathbf{x}_{\mathrm{hard}})=\epsilon$)
@@
    Intermediate ($\mathbf{x}_{\mathrm{inter}},~\epsilon\leq p(\mathbf{x}_{\mathrm{inter}})\leq 1-\epsilon$)
@@
    Soft-easy ($\mathbf{x}_{\mathrm{easy}},~p(\mathbf{x}_{\mathrm{easy}})=1-\epsilon$)
@@
    Absolute-easy ($\mathbf{x}_{\mathrm{Easy}},~p(\mathbf{x}_{\mathrm{Easy}})=1$)
@@
where $\epsilon$ is a small positive constant satisfying $0\ll\epsilon<0.5$. The variance is zero if and only if $p(\mathbf{x})=0$ or $p(\mathbf{x})=1$, corresponding to absolute hard and absolute easy prompts, respectively.
3.2 Prompt-level learnability: theoretical analysis
We analyze the learnability of prompts under the soft reinforcement learning formulation defined in §3.1. Let the binary verifiable reward $r_{\mathrm{acc}}(x,y)\in\{0,1\}$ follow a Bernoulli distribution,
$p(x):=\mathbb{E}_{y\sim\pi_{\mathrm{init}}(\cdot|x)}[r_{\mathrm{acc}}(x,y)],$ (12)
which we refer to as the pass rate of prompt $x$. Given the definitions of the value function and log-ratio, the expected log ratio between $\pi_{\mathrm{init}}$ and the soft-optimal policy $\pi^{*}$ can be written as